Grade 12Mechanics

A particle of mass m moving with a speed v, hits elastically another stationary particle of mass 2m on a smooth horizontal circular tube of radius r. The time in which next collision takes place is?

Aoishi
8 Years agoGrade 12

1 Answer

Rituraj Tiwari
5 Years ago

To solve the problem regarding the collision of two particles on a smooth circular tube, we need to analyze the situation step by step, taking into account the principles of conservation of momentum and the nature of elastic collisions. Let's break it down systematically.

Understanding the Collision Dynamics

We have two particles: one with mass m moving with speed v and another stationary particle with mass 2m. When they collide elastically, both momentum and kinetic energy are conserved. Since the second particle is stationary, we can use this information to derive the velocities after the collision.

Step 1: Applying Conservation of Momentum

The total momentum before the collision must equal the total momentum after the collision. Before the collision, the momentum is:

  • Momentum of particle 1 (mass m): mv
  • Momentum of particle 2 (mass 2m): 0 (since it is stationary)

The total initial momentum is:

Initial Momentum = mv

Let’s denote the final velocities after the collision as v1 for particle 1 and v2 for particle 2. The conservation of momentum gives us:

mv = mv1 + 2mv2

Step 2: Applying Conservation of Kinetic Energy

For an elastic collision, we also need to conserve kinetic energy. The total kinetic energy before the collision is:

Initial Kinetic Energy = (1/2)mv2

After the collision, the kinetic energy is:

Final Kinetic Energy = (1/2)mv12 + (1/2)(2m)v22

Setting these equal gives us:

(1/2)mv2 = (1/2)mv12 + mv22

Step 3: Solving the Equations

From the momentum equation, we can express v1 in terms of v2:

v1 = v - 2v2

Substituting this into the kinetic energy equation leads to:

(1/2)m(v2) = (1/2)m(v - 2v2)2 + mv22

After simplification, we can find the expressions for v1 and v2. The typical outcome of such a collision would be:

  • Particle 1 (mass m) will reverse its direction and have a velocity of -v/3.
  • Particle 2 (mass 2m) will take off with a velocity of v/3.

Calculating the Time Until the Next Collision

Now that we have the velocities, we need to find the time until the next collision occurs. The two particles will move along the circular path of radius r. The time taken for a particle to traverse the circumference of the circular tube can be calculated using the formula:

Time = Distance / Speed

The circumference of the circle is:

C = 2πr

For particle 2 (mass 2m), moving at speed v/3:

Time2 = (2πr) / (v/3) = (6πr) / v

Since particle 1 is moving in the opposite direction with speed v/3, it will also take the same amount of time to reach the collision point again. Thus, the time until the next collision occurs is:

Time until next collision = (6πr) / v

Final Thoughts

The key takeaway from this problem is the application of conservation laws in elastic collisions and the understanding of motion along a circular path. By analyzing the situation using momentum and energy principles, we can effectively determine the dynamics of the particles and predict future interactions.